# Comparing Bayesian and Frequentist Inference in Biological Models: A Comparative Analysis of Accuracy, Uncertainty, and Identifiability

### v1

@@ 5.2 Results
LV-1 $x(t),y(t)$ $\alpha,\beta,\delta,\gamma,x(t),y(t)$ $\alpha,\beta,\delta,\gamma,x(0),y(0)$
LV-2 $y(t)$ only $\alpha,\beta,\gamma,y(t)$ $\delta,x(t)$ $\alpha,\beta,\delta,\gamma,x(0),y(0)$
LV-3 $x(t)$ only $\alpha,\delta,\gamma,x(t)$ $\beta,y(t)$ $\alpha,\beta,\delta,\gamma,x(0),y(0)$
@@
Table 2: Structural identifiability results for all models under different observation schemes. The table indicates which parameters and state variables are identifiable or unidentifiable when initial conditions (ICs) are known versus unknown. LV-1, LV-2, and LV-3 refer to Lotka–Volterra scenarios with both species observed, predator only, and prey only, respectively.
@@ 5.4 Implications for Inference
The structural identifiability analysis provides critical context for interpreting the empirical parameter estimates and forecasting results presented in the Results section. For the LV model, we expect both Bayesian and Frequentist methods to successfully recover all parameters when both species are observed (LV-1), but parameter estimates may be unreliable or exhibit high uncertainty in the partial observation scenarios (LV-2, LV-3) where structural identifiability is compromised.
@@
Lotka–Volterra (LV) Ecology Hudson–Bay lynx–hare population Full (predator + prey), Prey-only, Predator-only
Generalized Logistic Model (GLM) Clinical Lung injury progression Full data
Generalized Logistic Model (GLM) Epidemiology 2022 U.S. mpox outbreak Full data
@@ 6 Results
In this section, we present parameter estimation results, performance metrics, and convergence diagnostics for all models and datasets analyzed in this study. We organize results by model, with each subsection containing parameter estimates, forecasting performance, and uncertainty quantification for both Bayesian (BFF) and Frequentist (QDF) inference methods. The structural identifiability analysis (Section 5) provides theoretical context for interpreting these empirical results.
@@ 6.1 Lotka–Volterra Model
We fitted the LV model to Hudson Bay lynx-hare data under three observation scenarios: (1) both prey and predator observed (LV-1), (2) predator only (LV-2), and (3) prey only (LV-3). These scenarios allow us to assess how partial observability affects parameter recovery and forecasting performance in light of the structural identifiability results presented in Section 5. To enhance the analysis, we consider these three distinct scenarios. First, we present the results of fitting the model to both prey and predator data simultaneously. Next, we focus on the scenario where the model is fitted exclusively to predator data. Finally, we provide the results for the case where only the prey data is used to fit the model.
6.1.1 Both Prey and Predator
In this section, we present the results with a focus on fitting the model to both prey and predator data.
Figure (5) shows that the BFF method is more effective in capturing the dynamics of the data. in particular, the prey population is modeled with a better range of uncertainty, covering all observed data points when BFF is applied. Furthermore, the method excels at capturing the peak of the predator population, particularly the first peak in the predator population, achieving a near-perfect fit.
Table 4), presents significantly different parameter estimates, despite using both datasets simultaneously. in particular, the parameter $\beta$ is estimated to be much lower in BFF compared to QDF, with the value approaching zero. Additionally, the parameter $\delta$ is the only one that shows approximately the same estimate in both methods.
Table 5) confirms the claim that BFF performs much better than QDF.
Table 6) presents the parameter $\hat{R}$ for all the parameters of the model, which is consistently 1.01. MCMC sampling was conducted with 2000 iterations, and these values of $\hat{R}$ ensure that the chains have converged successfully.
@@ 6.1.1 Both Prey and Predator
BFF .40 (.36, .44) .02 (.02, .02) .03 (.03, .03) 1.06 (.95, 1.17)
QDF .55 (.51,.60) .30 (.32,.35) .03 (.02,.03) .84 (.76,.92)
Table 4: Parameter estimates (joint predator–prey). $\beta$ differs across methods, while $\delta$ agrees.
The Figure 5 shows fitting the Hudson-Bay dataset to the LV model, assuming both prey and predator data simultaneously. The first row represents the results obtained using the BFF method, while the second row shows the results from the QDF method. The first column represents the predator population, while the second column corresponds to the prey population.
Figure 5: LV fit with joint predator–prey observations. Using both time series captures the coupled oscillations and aligns peaks and troughs in each species, indicating a well-calibrated interaction structure.
@@
Table 5: Forecast metrics (joint predator–prey). QDF lowers errors for predator; BFF attains perfect prey coverage.
@@
$\hat{R}$ (BFF) 1.01 1.01 1.01 1.01
Table 6: Convergence diagnostics (BFF). All $\hat{R}\approx 1.01$.
In the next sections, We are comparing the performance of two different fitting approaches: BFF and QDF. Since we apply both methods separately when fitting only one component of the dataset (either predator or prey), we should expect to generate two figures, one for each method’s fit.
6.1.2 Predator data only
In the following, we present the results of fitting the Hudson-Bay dataset to the LV model, considering only predator data.
@@ 6.1.2 Predator data only
Predator BFF 1.2(.75,1.9) .03(.01,.06) .02(.01,.03) .36(.22,.62)
QDF .97(.64,1) .02(.02,.06) .025(.02,.036) .46(.4,.7)
Table 7: The parameter estimation while fitting the Hudson-Bay dataset to the LV model, assuming predator data only. The first row represents the results obtained using the BFF method, while the second row shows the results from the QDF method.
Figure 6 below depicts fitting the Hudson-Bay dataset to the Lotka-Volterra model, assuming only predator. The first column represents the results obtained using the BFF method, while the second column shows the results from the QDF
Figure 6: Predator-only fit under the LV model. Calibrating to a single species reproduces coarse predator cycles but leaves prey dynamics unconstrained, reflecting weaker identifiability of interaction terms.
Table 8 below presents the performance metrics, showing that QDF outperforms BFF in terms of MAE, MSE, and WIS, indicating lower error and better predictive performance.
@@
Predator BFF 3.6 18.07 2.22 100.00
QDF 2.8 19.7 2.33 95.24
Table 8: The performance metrics while fitting the Hudson-Bay dataset to the LV model, assuming predator data only.
@@
Predator $\hat{R}$ (BFF) 1 1 1 1
Table 9: The convergence analysis while fitting the Hudson-Bay dataset to the LV model, assuming predator data only.
The results indicate that QDF provides a better fit to predator data alone, capturing key trends in the population more effectively than BFF.
6.1.3 Prey data only
In the following, we present the results of fitting the Hudson-Bay dataset to the LV model, considering only the prey population.
@@ 6.1.3 Prey data only
Prey BFF .34(.28,.52) .01(0,.05) .04(.02,.05) 1.4(.75,1.9)
QDF .42(.34,.55) .013(.005,.035) .03(.02,.04) 1.2(.78,1.5)
Table 10: The parameter estimation while fitting the Hudson-Bay dataset to the Lotka-Volterra model, assuming prey data only . The first row represents the results obtained using the BFF method, while the second row shows the results from the QDF method.
Figure 7 below illustrate fitting the Hudson-Bay dataset to the Lotka-Volterra model, assuming prey data only. The first column represents the results obtained using the BFF method for prey data only, while the second column shows the results from the QDF method for prey data only.
Figure 7: Prey-only fit under the LV model. Fitting to prey alone matches overall prey amplitude but yields ambiguous predator dynamics, again indicating partial identifiability with single-species data.
@@
Table 11: The performance metrics while fitting the Hudson-Bay dataset to the LV model, assuming prey data only.
@@
Prey $\hat{R}$ (BFF) 1.01 1.01 1.01 1.01
Table 12: The convergence analysis while fitting the Hudson-Bay dataset to the LV model, assuming prey data only.
These results suggest that QDF provides a more robust and accurate model for prey population dynamics, outperforming BFF in terms of both error metrics and parameter estimates.
6.1.4 Comparative analysis of parameter estimates ($\alpha,\beta,\delta,\gamma$) for the LV predator-prey model across observation scenarios (prey-only, predator-only, and predator-prey)
Figure 8 below presents a comparative analysis of the parameter estimates ($\alpha,\beta,\delta,\gamma$) for the (LV) predator-prey model. The estimates are shown in three observation scenarios: prey-only, predator-only, and predator-prey. Each point represents the mean estimate, and the horizontal bars denote the associated uncertainty intervals. Logarithmic scaling is applied to better visualize variations across small-magnitude parameter ranges. This comparison highlights how the availability of data influences the identifiability of the parameter and the consistency between the forecasting methodologies.
Figure 8: Parameter comparison across observation schemes. Joint data produce tighter intervals; single-series widens uncertainty.
6.1.5 Performance metrics comparison of MAE, MSE, WIS, and 95% PI for predator-only vs. predator–prey data in the LV model
Figure 9 below evaluates the predictive performance of the BFF and QDF frameworks under two data availability scenarios: (1) when only predator data are observed and (2) when predator and prey data are available. Forecast accuracy is assessed using four metrics: Mean Absolute Error (MAE), Mean Squared Error (MSE), Weighted Interval Score (WIS) and 95% PI. Across all metrics, using full predator-prey data improves model performance compared to relying on predator-only data, underscoring the value of joint observations in dynamic ecological systems.
@@
6.1.6 Performance metrics comparison of MAE, MSE, WIS, and 95% PI for prey-only vs. predator–prey data in the LV model
Figure 10 below evaluates the predictive performance of the BFF and QDF frameworks under two data availability scenarios: (1) when only prey data are observed and (2) when both predator and prey data are available. Forecast accuracy is assessed using four metrics: Mean Absolute Error (MAE), Mean Squared Error (MSE), Weighted Interval Score (WIS) and 95% PI. Across all metrics, using full predator-prey data improves model performance compared to relying on prey-only data, underscoring the value of joint observations in dynamic ecological systems.
@@
6.1.7 Posterior Histograms of LV Model Parameters $(\alpha,\beta,\delta,\gamma)$ Across Data Scenarios
The posterior histograms of the LV parameters $(\alpha,\beta,\delta,\gamma)$ are shown in Figures (S1)–(S6). These plots compare predator–prey, predator-only and prey-only scenarios under both BFF and QDF. Each histogram summarizes parameter uncertainty with mean, median and 95% confidence interval (CI), providing insight into identifiability and precision across data scenarios.
Specifically, Figures (S1)–(S3) show the BFF results, while Figures (S4)–(S6) present the QDF estimates. Together, they highlight method-specific differences while confirming the model’s ability to recover biologically meaningful values. Narrow, sharply peaked histograms (e.g., $\alpha\approx 0.4\pm 0.04$ in (Figure S1) indicate strong identifiability and low posterior uncertainty. Broad or flat histograms (e.g., predator-only $\beta$ or $\gamma$ in (Figure S2) reflect weak identifiability, these parameters can vary widely while still fitting the data. Comparing BFF vs QDF panels shows that both recover biologically reasonable values, but QDF tends to yield slightly tighter intervals when data coverage is richer.
6.1.8 Comparison of the parameters of the LV model (BFF) using a Normal distribution as the error structure.
@@ 6.1.7 Posterior Histograms of LV Model Parameters $(\alpha,\beta,\delta,\gamma)$ Across Data Scenarios
Table 13: Prior distributions and bounds for the parameters in the LV model. The priors are truncated normal distributions with mean 0.5 and standard deviation 0.5, constrained to be non-negative (T[0,]).
6.1.9 Comparison of the parameters of the LV model (QDF) using a Normal distribution as the error structure.
@@ 6.2 GLM model for Lung Injury
This section summarizes the results of fitting the generalized logistic model (GLM) to the lung-injury time series. We compare model fits, parameter estimates, predictive accuracy (MAE, MSE, WIS), and the coverage of 95% PI. Figure 11 shows that both methods capture the rise and fall of the affected population and align well at peaks and during the decline phase. Intervals are tight around turning points, with small differences between methods.
Figure 11: GLM fits to the lung-injury data using frequentist inference (QDF) and Bayesian inference (BFF); both capture the overall rise–fall pattern.
Table 15 reports $r,p,$ and $k$ estimates; both methods agree closely on central values. The first row represents the results obtained using the BFF method, while the second row shows the results from the QDF method.
@@
Table 15: Parameter estimates for the GLM (lung injury): both frameworks produce similar central estimates for $r$, $p$, and $k$.
Table 16 compares MAE, MSE, WIS, and 95% PI, where the frequentist QDF fit achieve lower errors with comparable coverage.
@@
Table 16: Performance metrics for the GLM (lung injury): QDF yields lower MAE, MSE, and WIS with comparable 95% PI.
Table 17 summarizes the convergence of the Bayesian fit (all $\hat{R}=1$), indicating good mixing.
@@
$\hat{R}$ (BFF) 1 1 1
Table 17: Convergence (Bayesian GLM, lung injury): all parameters have $\hat{R}=1$, indicating good MCMC convergence.
6.2.1 Comparison of Parameter Estimates for the GLM Model using Lung Injury data
Figure 12 shows the comparison of $r,p$ and $k$ between the methods, showing similar point estimates and slightly different uncertainty widths. Each point represents the posterior mean, with horizontal lines indicating the 95% PI. The comparison highlights the consistency in parameter estimation across methods while also revealing subtle differences in uncertainty ranges.
Figure 12: Parameter comparison for the GLM (lung injury): $r$, $p$, and $k$ agree across methods, with small differences in interval widths.
6.2.2 Performance metrics comparison of MAE, MSE, WIS, and 95% PI for the GLM Model using lung injury data
Figure 13 visualizes MAE, MSE, WIS, and PI . The results show that QDF consistently achieves lower error rates while maintaining comparable PI coverage, indicating more accurate and reliable forecasts.
Figure 13: Performance comparison for the GLM (lung injury): QDF attains lower errors while preserving PI.
6.2.3 Posterior Histograms of GLM Parameters ($r,k,p$) Across Data Scenarios
The posterior (or sampling) histograms of the GLM parameters $(r,k,p)$ are shown in Figures (S7)–(S8). Each set summarizes the uncertainty of parameter under BFF and QDF, reporting the mean, median, and 95% confidence interval (CI) for each parameter. Together, these panels allow for side-by-side comparison of how the two estimation frameworks explore $(r,k,p)$ across the data scenarios. Both BFF and QDF yield concentrated posteriors, for example, $p\approx 1.0\pm 0.02$ showing that these parameters are well identified.
6.2.4 Comparison of GLM Parameter Settings (BFF) with Normal Error
@@
Table 18: Parameter ranges for the GLM model.
6.2.5 Comparison of GLM Parameter Settings (QDF) with Normal Error

### v2

@@ 5.2 Results
Table 2: Structural identifiability results for all models under different observation schemes. The table indicates which parameters and state variables are identifiable or unidentifiable when initial conditions (ICs) are known versus unknown. LV-1, LV-2, and LV-3 refer to Lotka–Volterra scenarios with both species observed, predator only, and prey only, respectively. “NULL” indicates that no states or parameters are unidentifiable.
@@
LV-1 $x(t),y(t)$ $\alpha,\beta,\delta,\gamma,x(t),y(t)$ All states and parameters
LV-2 $y(t)$ only $\alpha,\beta,\gamma,y(t)$ $\delta,x(t)$ All states and parameters
LV-3 $x(t)$ only $\alpha,\delta,\gamma,x(t)$ $\beta,y(t)$ All states and parameters
@@ 5.4 Implications for Inference
The structural identifiability analysis provides critical context for interpreting the empirical parameter estimates and forecasting results presented in the Result section. For the Lotka–Volterra model, we expect both Bayesian and Frequentist methods to successfully recover all parameters when both species are observed (LV-1), but parameter estimates may be unreliable or exhibit high uncertainty in the partial observation scenarios (LV-2, LV-3) where structural identifiability is compromised.
@@
Lotka–Volterra Ecology Hudson–Bay lynx–hare population (predator + prey), Prey-only, Predator-only
GLM Clinical Lung injury progression Full data
GLM Epidemiology 2022 U.S. mpox outbreak Full data
@@ 6 Results
This section reports parameter estimates, calibration performance metrics, fitting visualization, and Bayesian convergence diagnostics across the models. We organize results by model and, when applicable, by observation scenario.
@@ 6.1 Lotka–Volterra Model
We fitted the Lotka–Volterra model to Hudson Bay lynx-hare data under three observation scenarios: (1) both prey and predator observed (LV-1), (2) predator only (LV-2), and (3) prey only (LV-3). These scenarios allow us to assess how partial observability affects parameter recovery and forecasting performance in light of the structural identifiability results presented in Section 5.
6.1.1 LV-1: Joint predator–prey observations
As shown in Figure 5, with both time series available, both methods recover the characteristic coupled oscillations with well-calibrated uncertainty. However, BFF provides better coverage of prey data between 1905-1910. Parameter estimates are almost similar for $\beta$ (0.02 vs. 0.025) and $\delta$ (both 0.03) but differ for other parameters, as shown in Table 4.
The calibration results show a trade-off that depends on the series (Table 5). For the predator series, QDF has lower errors, such as lower MAE and MSE. For the prey series, BFF performs better, with lower errors and higher coverage (100% compared to 90.48%). Table 6 also shows that the Bayesian chains mixed well, with $\hat{R}$ values close to 1 for all parameters.
Table 4: Parameter estimates for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing both predator and prey, using QDF and BFF approaches.
@@ 6.1.1 LV-1: Joint predator–prey observations
BFF 0.40 (0.36, 0.44) 0.02 (0.02, 0.02) 0.03 (0.03, 0.03) 1.06 (0.95, 1.17)
QDF 0.55 (0.51, 0.60) 0.025 (0.022, 0.029) 0.03 (0.02, 0.03) 0.84 (0.76, 0.92)
Figure 5: Fitting visualization for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing both predator and prey, using QDF and BFF approaches. Top row: BFF. Bottom row: QDF. Left column: predator. Right column: prey.
Table 5: Performance metrics for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing both predator and prey, using QDF and BFF approaches.
@@
Table 6: Convergence diagnostic for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing both predator and prey in Bayesian approach.
@@
$\hat{R}$ 1.01 1.01 1.01 1.01
6.1.2 LV-2: Predator-only observations
We next fit the Lotka–Volterra model using predator observations only. As expected under reduced observability (Section 5), parameter uncertainty increases and multiple parameter combinations can explain the observed predator cycles (Table 7). Despite this, point-fit performance is similar across methods (Table 8): QDF has slightly lower MAE (3.8 vs. 3.6 for BFF), while BFF has slightly lower MSE and WIS. PI coverage remains high for both (100% for BFF; 95.24% for QDF), and the Bayesian chains show $\hat{R}=1$ in this fit (Table 9).
Table 7: Parameter estimates for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing only predator, using QDF and BFF approaches.
@@ 6.1.2 LV-2: Predator-only observations
Predator BFF 1.20 (0.75, 1.90) 0.03 (0.01, 0.06) 0.02 (0.01, 0.03) 0.36 (0.22, 0.62)
QDF 0.97 (0.64, 1.00) 0.02 (0.02, 0.06) 0.025 (0.02, 0.036) 0.46 (0.40, 0.70)
BFF and QDF yield comparable predator-only fits (Figure 6), but implied prey dynamics are necessarily weakly constrained without prey observations.
Figure 6: Fitting visualization for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing only predator, using QDF and BFF approaches. Left BFF, and Right QDF.
Table 8: Performance metrics for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing only predator, using QDF and BFF approaches.
@@
Predator BFF 3.6 18.07 2.22 100
QDF 3.8 19.7 2.33 95.24
Table 9: Convergence diagnostic for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing only predator in Bayesian approach.
@@
Predator $\hat{R}$ 1 1 1 1
6.1.3 LV-3: Prey-only observations
Finally, we fit the Lotka–Volterra model using prey observations only. In this scenario, QDF achieves lower point-error metrics (Table 11: MAE 3.3 vs. 4.6; MSE 15.83 vs. 27.34; WIS 2.1 vs. 2.8), while BFF attains higher PI coverage (100% vs. 90.48%). Parameter estimates remain plausible but are less constrained than in the joint case (Table 10), and the Bayesian fit again shows $\hat{R}\approx 1.01$ (Table 12).
Table 10: Parameter estimates for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing only prey, using QDF and BFF approaches.
@@ 6.1.3 LV-3: Prey-only observations
Prey BFF 0.34 (0.28, 0.52) 0.01 (0.00, 0.05) 0.04 (0.02, 0.05) 1.40 (0.75, 1.90)
QDF 0.42 (0.34, 0.55) 0.013 (0.005, 0.035) 0.03 (0.02, 0.04) 1.20 (0.78, 1.50)
Both methods track the main prey fluctuations under prey-only calibration (Figure 7), but implied predator dynamics remain uncertain without direct predator observations.
Figure 7: Fitting visualization for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing only prey, using QDF and BFF approaches. Left BFF, and Right QDF.
@@
Table 11: Performance metrics for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing only prey, using QDF and BFF approaches.
@@
Table 12: Convergence diagnostic for the Lotka–Volterra model, observing only prey in Bayesian approach.
@@
Prey $\hat{R}$ 1.01 1.01 1.01 1.01
Across the three observation scenarios, parameter estimates and interval widths vary in a parameter-specific manner (Figure 8). Joint predator–prey observation primarily affects interaction-related parameters, most notably $\beta$, for which point estimates differ substantially across inference methods, reflecting stronger parameter coupling and improved identifiability. By contrast, for other parameters ($\alpha$, $\delta$, and $\gamma$), interval widths and estimates overlap considerably across observation scenarios, indicating that joint observation does not uniformly reduce uncertainty relative to single-series calibration.
Figure 8: Parameter comparison for the Lotka–Volterra model, with Hudson Bay lynx-hare data, across observation schemes: joint predator-prey, only predator, only prey. We use a log scale for the x-axis.
Comparing predator-only vs. joint data availability, adding prey observations improves calibration performance for both methods across MAE/MSE/WIS and PI coverage (Figure 9), underscoring the value of joint observations for dynamic ecological systems.
@@
Likewise, adding predator observations to a prey-only fit improves performance across metrics for both frameworks (Figure 10).
@@
The histogram summaries (Figures 2025) reinforce these patterns: joint data produce tighter, more concentrated uncertainty summaries, while predator-only and prey-only settings yield broader distributions, consistent with reduced information under partial observation (Section 5).
Table 13: Prior distributions and bounds for the parameters in the Lotka–Volterra model. The priors are truncated normal distributions with mean 0.5 and standard deviation 0.5, constrained to be non-negative (T[0,]).
@@
Table 14: Parameter ranges for the Lotka–Volterra model of (QDF).
@@ 6.2 GLM model for Lung Injury
The GLM provides a well-identified baseline example in which the data support stable parameter recovery under both frameworks (Section 5). Both approaches capture the rise and decline in the lung-injury time series with QDF providing a better fit (Figure 11).
Figure 11: Fitting visualization for the GLM model, with lung-injury data, using QDF and BFF approaches. Left QDF, and Right BFF.
The inferred growth parameters are close across methods (Table 15), with $r$ near 0.32–0.33 and $p$ near 1.0, indicating consistent recovery of the growth shape. The main discrepancy is in the estimated carrying capacity $k$, where BFF yields a higher central estimate than QDF.
In calibration performance, QDF achieves lower point-error metrics (MAE 9.42 vs. 11.85; MSE 152.80 vs. 412.29; WIS 5.75 vs. 7.15) while both methods attain high PI coverage (Table 16). This pattern is consistent with an identifiable, low-dimensional setting where optimization-based estimation can yield very sharp point fits without compromising interval calibration.
The Bayesian fit converged well for this dataset ($\hat{R}=1$ for all GLM parameters; Table 17), supporting reliability of the reported posterior summaries.
Table 15: Parameter estimates for the GLM model, with lung-injury data, using QDF and BFF approaches.
@@
Table 16: Performance metrics for the GLM model, with lung-injury data, using QDF and BFF approaches.
@@
Table 17: Convergence diagnostics for the GLM model, with lung-injury data, using QDF and BFF approaches.
@@
$\hat{R}$ 1 1 1
The cross-method parameter comparison reinforces that $(r,p)$ are tightly constrained in both frameworks, with only modest differences in uncertainty widths (Figure 12).
Figure 12: Parameter comparison for the GLM model, with lung-injury data, using QDF and BFF approaches.
Across the summary metrics, the same ranking is apparent: QDF yields lower MAE/MSE/WIS while PI coverage remains comparable (Figure 13).
Figure 13: Performance metrics comparison for the GLM model, with lung-injury data, using QDF and BFF approaches.
The parameter bounds used for applying the BFF and QDF for this model is provided in Table 18 and 19.
Table 18: Parameter bounds for the GLM model, with lung-injury data, using QDF and BFF approaches.
@@
Table 19: Parameter bounds for the GLM model, with lung-injury data, using QDF and BFF approaches.